The sum of 2 whole number is 45 and their difference is less than 19.the number of all possible pairs is
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

Let's call the numbers x and y, with x being the bigger of the two:
x+y = 45; y = 45-x
x-y =< 19
x-(45-x) =< 19
x-(-x+45) =<19
2x-45 =<19
2x =< 64
x=< 32
y=> 45-32 = 13
So your numbers, in (x,y) form, are:
(32,13), (31,14), (30,15), (29,16), (28,17) and so on, until you get to (23,22)

Answer by MathTherapy(10552)

The sum of 2 whole number is 45 and their difference is less than 19.the number of all possible pairs is
The other person's answer is WRONG!
Correct answer: